# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 2.2.1. Log canonical thresholds
where $s_{D}$ is a (multivalued) holomorphic section cutting out $D$ and $||\cdot||_{D}$, $||\cdot||_{\Delta}$ denotes any fixed smooth metrics on the line bundles corresponding to $D$ and $\Delta$, respectively. In order to calculate the log canonical thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ we will instead use the well-known valuative definition of $\mathrm{lct}(X,\Delta;D)$ (see, for example, [25, Thm B.5]):
(2.5) $\mathrm{lct}(X,\Delta;D)=\inf_{v}\mathrm{lct}(X,\Delta;D)_{v},\,\,\,\,\mathrm{lct}_{v}(X,\Delta;D)\coloneqq\frac{A_{(X,\Delta)}(v)}{v(D)},$
@@
(2.6) $A_{(X,\Delta)}(v)=1+\mathrm{ord}_{E}(\pi^{*}K_{(X,D)}-K_{Y}),$
where $\pi^{*}K_{(X,D)}-K_{Y}$ defines a divisor, supported on the exceptional divisors of $\pi.$
@@
Let $\pi\colon Y\rightarrow X$ be a log resolution of $(X,D+\Delta),$ i.e. a birational morphism that is an isomorphism over the complement of $\text{supp}(D+\Delta)$ such that $\pi^{*}(D+\Delta)$ is a divisor on $Y$ with simple normal crossings. Denoting by $E_{1},\dots,E_{m}$ its components, the inf in 2.5 can be restricted to the divisorial valuations defined by $E_{1},\dots,E_{m}$. Using the existence of log resolutions it then follows readily that the two definitions 2.4 and 2.5 of the log canonical threshold agree.
@@
Next, if $U$ is an open subset in the Euclidean topology of $X$, then $\mathrm{lct}(U,\Delta|_{U},D|_{U})$ designates the inf 2.5 restricted to valuations whose center have non-zero intersection with $U$. With this definition, analogous versions of the above properties hold, and moreover, if $X=\cup_{i}U_{i}$, then clearly,
@@
A divisorial valuation $v$ over $X$ is said to compute $\mathrm{lct}(X,\Delta;D)$ if the inf 2.5 is attained at $v.$ More generally, when considering a sequence $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ a sequence $v_{N}$ of valuations over $X_{N}$ will be said to asymptotically compute $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ if $\text{lct}_{v_{N}}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\liminf_{N\rightarrow\infty}\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),$ as $N\rightarrow\infty.$ Finally, note that any $N$-tuple of divisorial valuations $v_{1},\dots,v_{N}$ over $X$ induces a valuation over $X^{N}$ by blowing up the variety $E_{1}\times\cdots\times E_{N}$ over $X^{N},$ where $E_{i}$ is the divisor over $X$ corresponding to $v_{i}.$ Such a divisorial valuation over $X^{N}$ will be called decomposable.
@@ 2.2.2. Semistability in GIT
Let $L\rightarrow Y$ be an ample line bundle over a complex projective manifold $Y$ and $G$ a complex reductive group acting on $(Y,L).$ The corresponding semistable locus (in sense of Geometric Invariant Theory) is the Zariski open subset $Y_{\text{ss}}$ of $Y$ consisting of all $y\in Y$ such that there exists a GG–invariant section $s_{k}\in H^{0}(Y,kL)$ for some positive integer $k$ such that $s_{k}(y)\neq 0$ [64]. When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathrm{PGL}(2,\mathbb{C}),$ acting diagonally, $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the the fraction $I/N$ of points $x_{i}$ coinciding satisfies $I/N\leq 1/2$, i.e.
@@
for all $x\in\mathbb{P}^{1}$ (see [85, Section 3]). When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathbb{C}^{*}$ the semistable locus $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the inequality 2.8 holds for the two fixed points $x$ of the action of $\mathbb{C}^{*}$ on $\mathbb{P}^{1}.$ Just like the case $G=\mathrm{PGL}(2,\mathbb{C}$) this follow readily from the Hilbert–Mumford criterion for one-parameter subgroups of $\mathbb{C}^{*}\subset G.$
@@ 2.2.3. The Fulton–MacPherson compactification of configuration spaces
    $X^{[N]}$ is a non-singular variety, projective if $X$ is [58, Theorems 1 and 2].
@@
    There is a canonical morphism $\pi_{\mathrm{FM}}\colon X^{[N]}\rightarrow X^{N}$, which restricts to an isomorphism $\pi_{\mathrm{FM}}\colon X^{[N]}\setminus Z\rightarrow X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\},$ where $Z$ defines a simple normal crossings divisor [58, Theorem 3].
@@
    where $\mathcal{I}(\cdot)$ denotes the corresponding ideal sheaves and the sum ranges over all $W^{\prime}$ in $\mathcal{W}$ contained in $E$ [58, Theorem 3].
@@
Indeed, $\pi_{W}^{*}(\mathcal{I}(W))=\mathcal{I}(E_{\pi_{W}})$ and thus by the commutativity of the diagram, $f^{*}(\mathcal{I}(E_{\pi_{W}}))=f^{*}\circ\pi_{W}^{*}(\mathcal{I}(W))=\pi_{FM}(\mathcal{I}(E_{\pi_{W}})),$ which, by property 3 of the Fulton–MacPherson compactification above, coincides with the ideal defined by the divisor $\sum_{W^{\prime}\subset W}E_{W^{\prime}}.$ Since the $E_{W}$’s are prime this proves the identity 2.9. It follows from 2.9 that
@@ 2.3.2. The case $\Delta\neq 0$
Since $\Delta_{N}|_{U}$ is trivial, we can use the Fulton–MacPherson compactification as a log resolution there, and we turn to the other cases. Without loss of generality we may assume that $q=(p_{i_{1}},\dots,p_{i_{1}},p_{i_{2}},\dots,p_{i_{2}},\dots,p_{i_{l}},\dots,p_{i_{l}})$ for $1\leq l\leq m$ and some $p_{i_{1}},\dots,p_{i_{l}}$ distinct and where $p_{i_{1}}$ is repeated $k_{1}$ times and $p_{i_{2}}$ is repeated $k_{2}$ times, and so on. Then $U_{q}=V_{1}\times\cdots\times V_{l}\times(\mathbb{P}^{1})^{r}$, for some $0\leq r\leq N-1$, where $V_{j}=\{(x_{1},\dots,x_{k_{j}})\in(\mathbb{P}^{1})^{k_{j}}\colon d(x_{j},p_{i_{j}})<d_{0}\forall j\}$. Observe that $\Delta_{N}|_{U_{q}}=\Delta_{k_{1}}^{p_{i_{1}},w_{i_{1}}}\boxplus\cdots\boxplus\Delta_{k_{l}}^{p_{i_{l}},w_{i_{l}}}\boxplus 0$. Furthermore, $\mathcal{D}^{(N)}|_{U_{q}}\eqqcolon\mathcal{D}_{k_{1}}\boxplus\cdots\boxplus\mathcal{D}_{k_{l}}\times\mathcal{D}_{r}$. By 2.7
@@
But for any $1\leq j\leq l$, $\pi^{*}(\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}})\leq\Delta_{N}^{p_{i_{j}},w_{i_{j}}}$, and $\pi^{*}(\mathcal{D}_{k_{j}})\leq\mathcal{D}^{(N)}$ where $\pi\colon(\mathbb{P}^{1})^{N}\rightarrow(\mathbb{P}^{1})^{k_{j}}$ is the projection and similarly $\pi^{*}(\mathcal{D}_{r})\leq\mathcal{D}^{(N)}$. In view of the definition 2.5 and 2.7, $\mathrm{lct}(V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}_{k_{j}})\geq\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ with $q^{\prime}=(p_{i_{j}},\dots,p_{i_{j}})$ but the numbers $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ are contenders in the inf 2.13. Moreover $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})=\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}^{(N)})$ and any such term is lesser or equal to $\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U})$ in 2.14, and also $\mathrm{lct}((\mathbb{P}^{1})^{r};\mathcal{D}_{r})$. This proves formula 2.12.
@@
Now regard these as functions of $N+1$ variables $z_{1},\dots,z_{N}$ and $z_{N+1}$, which are independent of $z_{N+1}.$ In other words, we consider their pullbacks to $\mathbb{C}^{N+1}$ under the natural projection from $\mathbb{C}^{N+1}$ to $\mathbb{C}^{N}.$ This does not affect the computation of the corresponding log canonical threshold, by 2.7. After making the affine change of variables $z_{i}=u_{i}-u_{N+1}$ for $i=1,\dots,N$ and $z_{N+1}=u_{N+1}$ the two expressions 2.15 become
@@
The formula for $\gamma^{(N)}(X,\Delta)$ in the previous proposition corrects the formula given in [18, Thm 4.5], which is only correct modulo an error term of order $O(1/N)$ (that comes from the discrepancy between the upper and the lower bound on $\gamma^{(N)}(X,\Delta)$ established by analytic means in [18]).

### v3

@@ 2.1.1. Microscopic stability thresholds and Gibbs polystability
It follows directly from the definitions that
(2.5) $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}\geq\gamma^{(N)}(X,\Delta)\geq\alpha(X,\Delta):=\inf_{D\sim-K_{(X,\Delta)}}\text{lct}\left(X,\Delta;D\right),$
where $D$ ranges over all effective $\mathbb{Q}$-divisors $D$ linearly equivalent to $K_{(X,\Delta)}$. The invariant $\alpha(X,\Delta)$ coincides with Tian’s alpha invariant of $(X,\Delta),$ which is strictly positive for any log Fano manifold $(X,\Delta).$
@@ 2.2.1. Log canonical thresholds
where $s_{D}$ is a (multivalued) holomorphic section cutting out $D$ and $||\cdot||_{D}$, $||\cdot||_{\Delta}$ denotes any fixed smooth metrics on the line bundles corresponding to $D$ and $\Delta$, respectively. In order to calculate the log canonical thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ we will instead use the well-known valuative definition of $\mathrm{lct}(X,\Delta;D)$ (see, for example, [27, Thm B.5]):
(2.7) $\mathrm{lct}(X,\Delta;D)=\inf_{v}\mathrm{lct}(X,\Delta;D)_{v},\,\,\,\,\mathrm{lct}(X,\Delta;D)_{v}\coloneqq\frac{A_{(X,\Delta)}(v)}{v(D)},$
@@
(2.8) $A_{(X,\Delta)}(v)=1+\mathrm{ord}_{E}(K_{Y}-\pi^{*}K_{(X,D)}),$
where $K_{Y}-\pi^{*}K_{(X,D)}$ defines a divisor, supported on the exceptional divisors of $\pi.$
@@
Let $\pi\colon Y\rightarrow X$ be a log resolution of $(X,D+\Delta),$ i.e. a birational morphism that is an isomorphism over the complement of $\text{supp}(D+\Delta)$ such that $\pi^{*}(D+\Delta)$ is a divisor on $Y$ with simple normal crossings. Denoting by $E_{1},\dots,E_{m}$ its components, the inf in 2.7 can be restricted to the divisorial valuations defined by $E_{1},\dots,E_{m}$. Using the existence of log resolutions it then follows readily that the two definitions 2.6 and 2.7 of the log canonical threshold agree.
@@
Next, if $U$ is an open subset in the Euclidean topology of $X$, then $\mathrm{lct}(U,\Delta|_{U},D|_{U})$ designates the inf 2.7 restricted to valuations whose center have non-zero intersection with $U$. With this definition, analogous versions of the above properties hold, and moreover, if $X=\cup_{i}U_{i}$, then clearly,
@@
A divisorial valuation $v$ over $X$ is said to compute $\mathrm{lct}(X,\Delta;D)$ if the inf 2.7 is attained at $v.$ More generally, when considering a sequence $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ a sequence $v_{N}$ of valuations over $X_{N}$ will be said to asymptotically compute $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ if $\text{lct}_{v_{N}}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\liminf_{N\rightarrow\infty}\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),$ as $N\rightarrow\infty.$ Finally, note that any $N$-tuple of divisorial valuations $v_{1},\dots,v_{N}$ over $X$ induces a valuation over $X^{N}$ by blowing up the variety $E_{1}\times\cdots\times E_{N}$ over $X^{N},$ where $E_{i}$ is the divisor over $X$ corresponding to $v_{i}.$ Such a divisorial valuation over $X^{N}$ will be called decomposable.
@@ 2.2.2. Semistability in GIT
Let $L\rightarrow Y$ be an ample line bundle over a complex projective manifold $Y$ and $G$ a complex reductive group acting on $(Y,L).$ The corresponding semistable locus (in sense of Geometric Invariant Theory) is the Zariski open subset $Y_{\text{ss}}$ of $Y$ consisting of all $y\in Y$ such that there exists a GG–invariant section $s_{k}\in H^{0}(Y,kL)$ for some positive integer $k$ such that $s_{k}(y)\neq 0$ [83]. More generally, given an ample line bundle $L\rightarrow Y$ with an action of $G$ on $(X,mL)$ for some positive integer $m,$ the semistable locus $Y_{\text{ss }}$ is defined wrt $(X,mL).$ When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathrm{PGL}(2,\mathbb{C}),$ acting diagonally, $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the the fraction $I/N$ of points $x_{i}$ coinciding satisfies $I/N\leq 1/2$, i.e.
@@
for all $x\in\mathbb{P}^{1}$ (see [103, Section 3]). When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathbb{C}^{*}$ the semistable locus $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the inequality 2.10 holds for the two fixed points $x$ of the action of $\mathbb{C}^{*}$ on $\mathbb{P}^{1}.$ Just like the case $G=\mathrm{PGL}(2,\mathbb{C}$) this follow readily from the Hilbert–Mumford criterion for one-parameter subgroups of $\mathbb{C}^{*}\subset G.$
@@ 2.2.3. The Fulton–MacPherson compactification of configuration spaces
    $X^{[N]}$ is a non-singular variety, projective if $X$ is [74, Theorems 1 and 2].
@@
    There is a canonical morphism $\pi_{\mathrm{FM}}\colon X^{[N]}\rightarrow X^{N}$, which restricts to an isomorphism $\pi_{\mathrm{FM}}\colon X^{[N]}\setminus Z\rightarrow X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\},$ where $Z$ defines a simple normal crossings divisor [74, Theorem 3].
@@
    where $\mathcal{I}(\cdot)$ denotes the corresponding ideal sheaves and the sum ranges over all $W^{\prime}$ in $\mathcal{W}$ contained in $E$ [74, Theorem 3].
@@
Indeed, $\pi_{W}^{*}(\mathcal{I}(W))=\mathcal{I}(E_{\pi_{W}})$ and thus by the commutativity of the diagram, $f^{*}(\mathcal{I}(E_{\pi_{W}}))=f^{*}\circ\pi_{W}^{*}(\mathcal{I}(W))=\pi_{FM}(\mathcal{I}(E_{\pi_{W}})),$ which, by property 3 of the Fulton–MacPherson compactification above, coincides with the ideal defined by the divisor $\sum_{W^{\prime}\subset W}E_{W^{\prime}}.$ Since the $E_{W}$’s are prime this proves the identity 2.11. It follows from 2.11 that
@@ 2.3.2. The case $\Delta\neq 0$
Since $\Delta_{N}|_{U}$ is trivial, we can use the Fulton–MacPherson compactification as a log resolution there, and we turn to the other cases. Without loss of generality we may assume that $q=(p_{i_{1}},\dots,p_{i_{1}},p_{i_{2}},\dots,p_{i_{2}},\dots,p_{i_{l}},\dots,p_{i_{l}})$ for $1\leq l\leq m$ and some $p_{i_{1}},\dots,p_{i_{l}}$ distinct and where $p_{i_{1}}$ is repeated $k_{1}$ times and $p_{i_{2}}$ is repeated $k_{2}$ times, and so on. Then $U_{q}=V_{1}\times\cdots\times V_{l}\times(\mathbb{P}^{1})^{r}$, for some $0\leq r\leq N-1$, where $V_{j}=\{(x_{1},\dots,x_{k_{j}})\in(\mathbb{P}^{1})^{k_{j}}\colon d(x_{j},p_{i_{j}})<d_{0}\forall j\}$. Observe that $\Delta_{N}|_{U_{q}}=\Delta_{k_{1}}^{p_{i_{1}},w_{i_{1}}}\boxplus\cdots\boxplus\Delta_{k_{l}}^{p_{i_{l}},w_{i_{l}}}\boxplus 0$. Furthermore, $\mathcal{D}^{(N)}|_{U_{q}}\eqqcolon\mathcal{D}_{k_{1}}\boxplus\cdots\boxplus\mathcal{D}_{k_{l}}\times\mathcal{D}_{r}$. By 2.9
@@
But for any $1\leq j\leq l$, $\pi^{*}(\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}})\leq\Delta_{N}^{p_{i_{j}},w_{i_{j}}}$, and $\pi^{*}(\mathcal{D}_{k_{j}})\leq\mathcal{D}^{(N)}$ where $\pi\colon(\mathbb{P}^{1})^{N}\rightarrow(\mathbb{P}^{1})^{k_{j}}$ is the projection and similarly $\pi^{*}(\mathcal{D}_{r})\leq\mathcal{D}^{(N)}$. In view of the definition 2.7 and 2.9, $\mathrm{lct}(V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}_{k_{j}})\geq\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ with $q^{\prime}=(p_{i_{j}},\dots,p_{i_{j}})$ but the numbers $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ are contenders in the inf 2.15. Moreover $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})=\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}^{(N)})$ and any such term is lesser or equal to $\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U})$ in 2.16, and also $\mathrm{lct}((\mathbb{P}^{1})^{r};\mathcal{D}_{r})$. This proves formula 2.14.
@@
Now regard these as functions of $N+1$ variables $z_{1},\dots,z_{N}$ and $z_{N+1}$, which are independent of $z_{N+1}.$ In other words, we consider their pullbacks to $\mathbb{C}^{N+1}$ under the natural projection from $\mathbb{C}^{N+1}$ to $\mathbb{C}^{N}.$ This does not affect the computation of the corresponding log canonical threshold, by 2.9. After making the affine change of variables $z_{i}=u_{i}-u_{N+1}$ for $i=1,\dots,N$ and $z_{N+1}=u_{N+1}$ the two expressions 2.17 become
@@
The formula for $\gamma^{(N)}(X,\Delta)$ in the previous proposition corrects the formula given in [19, Thm 4.5], which is only correct modulo an error term of order $O(1/N)$ (that comes from the discrepancy between the upper and the lower bound on $\gamma^{(N)}(X,\Delta)$ established by analytic means in [19]).